OpenStudy (anonymous):

A child is creating a pyramid with building blocks. The top three levels include 5 blocks, 12 blocks, and 19 blocks. Part 1: How many blocks would be needed for a pyramid 25 levels tall? (5 points) Part 2: Use complete sentences to explain how a sum of an arithmetic series was applied. (4 points)

OpenStudy (superdavesuper):

level 1: 5 blocks level 2: 12 = 5+7*1 level 3: 19=5+7*2 ..... level n: how many blocks u think are there?

OpenStudy (anonymous):

level 3: 26 level 4: 33 level 5: 40 level 6: 47 level 7: 54 level 8: 38 ? that correct going up?

OpenStudy (superdavesuper):

yes and the general eqn for n level is 5+7*(n-1) ok?

OpenStudy (anonymous):

yes and that would be for part 2 thank you

OpenStudy (superdavesuper):

welcome - what about part 1? :)

OpenStudy (anonymous):

i would just keep on going up

OpenStudy (superdavesuper):

thats one way of doing it....but theres an easier way....

OpenStudy (superdavesuper):

notice every level, there are 5+something blocks...so for 25 levels, it would be 25*(5)+something block right?

OpenStudy (anonymous):

ok 180 blocks

OpenStudy (superdavesuper):

try again pl? 25*5 is....

OpenStudy (anonymous):

125

OpenStudy (superdavesuper):

:) good...so now we have to add the +something part of each level...okay?

OpenStudy (anonymous):

+7?

OpenStudy (superdavesuper):

well +7 is part of it but it is +7*(n-1) for n level right?

OpenStudy (anonymous):

so +7*(25-1)?

OpenStudy (superdavesuper):

That is correct for the 25th level :) so all we need is to add from 1,2,3...to 25 levels to find the total...ok?

OpenStudy (superdavesuper):

Nope...do u understand everything up to this point first? :)

OpenStudy (anonymous):

yea

OpenStudy (superdavesuper):

OK so adding them together...7*0+7*1+7*2+.....7*(25-1), we can group them and it becomes 7*(0+1+2+....+24) okay?

OpenStudy (anonymous):

25*5+(12*24+12)*7 do this and i get the answer?

OpenStudy (superdavesuper):

hahahaha u already know the eqn for an arithmetic series? u hardly need my help :P

OpenStudy (superdavesuper):

n yes, ur ans is right - good job!

OpenStudy (anonymous):

2225 thats it

OpenStudy (superdavesuper):

u got it :)

OpenStudy (anonymous):

thanks

OpenStudy (superdavesuper):

i hate writing complete sentences....u good for part 2 i hope? ;)

OpenStudy (anonymous):

kind of

OpenStudy (superdavesuper):

cool - let me look at ur new Q....not sure i have enough time...